OpenStudy (anonymous):

Suppose that y varies directly as x and inversely as the square of z. When x = 3 and z = 5, y = 4.5. Find y when x = –2 and z = 3. A – 25/3 B 25/3 C –3/25 D 3/25

OpenStudy (anonymous1):

y varies directly as x and inversely as the square of z can be written as: \[y=k\frac{x}{z^2}\] where k is a constant.

OpenStudy (anonymous1):

"y varies directly as x" means that y is some constant k multiplied by x. "y varies inversely as the square of z" means that y is some constant divided by the square of z. Hence the equation I wrote above.

OpenStudy (anonymous1):

We have to find the value of the constant k. We know that y=4.5 then x=3 and z=5. Let's plug in these values: \[y=k\frac{x}{z^2}\] \[4.5=k\frac{3}{5^2}\] \[4.5=k\frac{3}{25}\] \[k=\frac{4.5\times25}{3}=1.5\times25\] k=37.5

OpenStudy (anonymous1):

Rewriting the equation with the value we discovered for k: \[y=37.5\frac{x}{z^2}\] Now we can find y when x=-2 and z=3: \[y=37.5\frac{-2}{3^2}\] \[y=37.5\frac{-2}{9}\] \[y=-8.3333...\] The fraction 8.333... is equivalent to 25/3. Therefore: \[y=-\frac{25}{3}\] So, the correct answer is A.